### v2

Abstract
This work is motivated by (i) the recently proposed relationship between gravitational tension (through the Kretschmann scalar) and energy density Estrada and Aros (2025) to construct regular black hole (RBH) solutions in Pure Lovelock (PL) gravity, and (ii) the fact that, for PL, the inclusion of a negative cosmological constant $\Lambda<0$ leads to the appearance of a potential curvature singularity Cai and Ohta (2006). Thus, by using a particular choice of coupling constants such that the resulting equations of motion for Lovelock gravity have an $n$-fold degenerate ground state (LnFDGS) AdS, we provide a methodology to construct RBH solutions with $\Lambda<0$ using an energy density model analogous to (but different in the definition of gravitational tension from) the one mentioned earlier. Moreover, because relating the gravitational tension to the Kretschmann scalar of the vacuum LnFDGS solution is complicated, we define an alternative version of the Kretschmann scalar suitable for the LnFDGS AdS theory and, consequently, a redefinition of the gravitational tension. Remarkably, we obtain a model where there exists a value $r_{*}$ slightly greater than the extremal radius, $r_{*}>r_{ext}$, which could be on the order of the Planck length, such that the solutions of the vacuum AdS black hole and our AdS RBH become indistinguishable. However, at short length scales such that $r<r_{*}$, quantum effects would arise, causing both cases to differ in their geometry (suppressing the central singularity) and their thermodynamic properties. Additionally, since it is not possible to find analytical relationships between the event horizon, the photon sphere radius, and the shadow size in LnFDGS, we propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior.
I Introduction
It is widely recognized that in the last decades various branches of theoretical physics predict the existence of extra dimensions. Although several experiments have attempted to test this hypothesis, it has not yet been observed. As a result, any theory that incorporates extra dimensions must be consistent with General Relativity in four dimensions. One such theory is Lovelock gravity. The Lagrangian of Lovelock gravity incorporates higher curvature terms as corrections to the Einstein-Hilbert action Lovelock (1971). Additionally, Lovelock’s theories respect the core principles of General Relativity; for instance, their equations of motion are of second order. It is worth noting that the specific case of Lovelock gravity known as Einstein-Gauss-Bonnet theory has gained attention in recent years for its applications in inflationary models and has been compared with the results from GW170817 Oikonomou (2021). See recent applications in references Bueno et al. (2025).
On the other hand, The singularity theorem, see Hawking and Ellis (2023); Penrose (1965), considered a triumph of theoretical physics, states in general terms that the collapse of matter could lead to the formation of singularities. However, it also suggests, broadly speaking, that the dynamics of gravity result in the breakdown of its own predictability. To address this fundamental issue, one could argue that the prediction of singularities arises from the incompleteness of current theories of gravity.
Although some theories that incorporate quantum effects, such as String Theory and Loop Quantum Gravity, to name just two examples, provide fundamentally different solutions to the singularity problem, it is not necessary to have a fully developed (fundamental) quantum gravity theory to offer solutions to the existence of singularities. This leads to ideas associated with the fact that gravity should be modified at small distances, typically around the Planck length. Therefore, in a sense, the removal or smoothing of these singularities should come from the incorporation of quantum effects in the matter fields present in the energy-momentum tensor. This latter idea has led to solutions known as Regular Black Holes (RBH), where the structure of the energy-momentum tensor is such that singularities at the origin are suppressed, generally leading to the formation of a core with a de Sitter structure rather than a singularity at the radial origin.
One way that has attracted attention in recent years to define the energy-momentum tensor, in order to address the existence of singularities, is by establishing a relationship between the gravitational tension of the vacuum case (associated with the Kretschmann scalar) and the energy density. It is well known that an invariant associated with the measurement of tidal forces is the Kretschmann scalar Bena et al. (2021). In a four-dimensional, vacuum, spherically symmetric space, this scalar is proportional to $K\sim{M}^{2}/r^{6}$, leading to infinite tidal forces at the origin. Gravitational field tension in the spherically symmetric case is characterized by the curvature term given by the square root of the Kretschmann scalar of the vacuum solution, $F\sim\sqrt{K_{\text{Schw}}}\sim\frac{{M}}{r^{3}}$ in $4D$. This correlation is logical, as the spacetime tension should increase with the mass of the vacuum source Alencar et al. (2023). In this conection, as a way to address the problem of singularities while including an appropriate structure for the energy-momentum tensor, a higher-dimensional model was recently constructed in reference Estrada and Aros (2025) for Pure Lovelock (PL) gravity, where the energy density encodes the gravitational information of the vacuum solution through the Kretschmann scalar. Near the radial origin, where tidal forces and the gravitational tension in the vacuum solution diverge, the tidal forces in the model become finite. Specifically, the energy density is written as:
$\rho\sim\exp\left(-\frac{F_{c}}{F}\right)$ (1)
being $F_{c},a$ constants. This model coincides with the RBH of Dymnikova Dymnikova (1992) in the four dimensional case. On the other hand, in a speculative sense, in four dimensions, this way of relating gravitational tension to energy density is analogous to the particle production ratio in the quantum Schwinger effect DYMNIKOVA (1996). We can find studies related to this latter topic in the references Ansoldi (2008); Alencar et al. (2023); Estrada and Muniz (2023) . However, this is a topic that requires deeper investigation.
On the other hand, over the last 50 years, much has been discussed regarding how General Relativity and other theories of gravity can be improved to address theoretical inconsistencies and mismatches with observations. Naturally, in recent decades, there has been a growing interest in determining the role of the cosmological constant in the functioning of gravity. Although the expansion of the universe has usually been associated with the existence of a positive cosmological constant, this fact brings some complications from a theoretical standpoint. Firstly, the presence of a positive cosmological constant is associated with the existence of a cosmological horizon, which prevents the existence of an asymptotic regime, making the unambiguous identification of the ADM mass of a black hole problematic. Secondly, a positive cosmological constant is usually associated with negative pressure, implying thermodynamic instability.
On the contrary, the presence of a negative cosmological constant offers good explanations for some physical ambiguities. A spacetime with a negative cosmological constant allows for the definition of an asymptotic boundary in the geometry of Penrose diagrams. In this context, the existence of an asymptotic boundary enables the correct definition of conserved charges Aros et al. (2001), such as mass/energy. The existence of a well-defined boundary in an AdS universe has led to the development of the AdS/CFT theory. In broad terms, the AdS/CFT theory is a correspondence between a quantum theory defined in a certain class of anti-de Sitter space and a conformal field theory defined on its boundary with one lower dimension. Moreover, it is worth mentioning that in recent years, attention has been drawn to associating the thermodynamics of an AdS Schwarzschild space, where the negative cosmological constant has been linked to a positive thermodynamic pressure Kubiznak and Mann (2012). This has allowed for the study of phase transitions, where stable regions can be found. In this way, it is of physical interest to study how the presence of a negative cosmological constant influences the thermodynamic properties of black holes.
A special case of Lovelock theories is Pure Lovelock theory (PL) Cai and Ohta (2006); Dadhich et al. (2012). The latter theory mentioned considers only a single term in the Lagrangian. Related to the aforementioned, a disadvantage of PL theory lies in the fact that considering the presence of a negative cosmological constant in a spherically symmetric spacetime leads to the appearance of a curvature singularity. This singularity is such that, although the metric tensor is regular for a radius greater than the event horizon, the Ricci and Kretschmann invariants diverge at a point within this location. This issue has not been extensively investigated, and thus, it may be related to a potential breakdown of spacetime. Therefore, studying the recent analogy between gravitational tension and energy density, as mentioned earlier, seems to be unfeasible in PL theory under the presence of a negative cosmological constant for the construction of RBH solutions.
One potential drawback of generic Lovelock gravity is the existence of multiple ground states, meaning more than one solution with constant curvature spaces, or equivalently, more than one potential effective cosmological constant Camanho and Edelstein (2013). These effective cosmological constants can be complex numbers, which can lead to instability of the ground states under dynamical evolution. To address this issue, one can choose the coupling constant in such a way that the equations of motion roughly take the form $(R-\Lambda)^{n}=0$ Crisostomo et al. (2000). In this case, there is a single ground state, but it is n-fold degenerate with constant curvature. This scenario is referred to as Lovelock with a unique vacuum (LUV) or Lovelock with an n-fold degenerate ground state (LnFDGS). In the context of this work, this coupling constant structure allows for the observation of a single negative effective cosmological constant, which, among other things, has facilitated the development of techniques for determining conserved charges Aros and Estrada (2019). This also seems to make sense in the framework of the AdS/CFT correspondence, as two or more negative cosmological constants would lead to different asymptotic boundary structures for the same background. See some recent applications of LnFDGS theory in references Xu and Yung (2019); Aros and Estrada (2019); Estrada and Aros (2019); Aros et al. (2022); Estrada (2022); Aros and Estrada (2024)
PL has significant differences compared to the LnFDGS theory Aros and Estrada (2019); Crisostomo et al. (2000). While PL considers only a single term from the Lovelock series, LnFDGS includes all the terms from this series up to $n=[d/2]$, where the square brackets represent the floor function. Moreover, PL cannot be factorized into several equivalent negative effective cosmological constants like LnFDGS. Remarkably, LnFDGS does not exhibit a curvature singularity of the type mentioned earlier. The latter offers the advantage of displaying a well-defined asymptotically AdS structure. In the context of this work, the absence of the aforementioned pathology in LnFDGS allows for the construction of RBH solutions with the presence of a negative cosmological constant, and for testing the physical consequences of its presence from both the geometric and thermodynamic perspectives.
On the other hand, it is well known that light is not directly observable at the event horizon of a black hole. Instead, what can be seen is the so-called shadow of a black hole, a dark region surrounded by light that results from the gravitational lensing phenomenon around the event horizon. In this way, an important achievement of the Event Horizon Telescope (EHT) collaboration has been the construction of images of the shadows of the supermassive black holes M87 Akiyama et al. (2021) and Sgr A Akiyama et al. (2022). In this regard, testing the behavior of shadows in scenarios that consider extra dimensions has also attracted attention: see, for example, reference Nozari et al. (2024) for Scalar-Tensor-Vector theories and reference Singh et al. (2023) (and the references therein) for rotating black hole shadows. It is important to mention that in reference Paithankar and Kolekar (2023), connections were established between the parameters of the black hole solution and the structure of the shadow for extra-dimensional and asymptotically flat scenarios. This study was carried out for scenarios where the event horizon radius can be determined analytically. Thus, relationships were established between the size of the latter and the radii of the photon sphere and the shadow. The vacuum solutions of Pure Lovelock and $f(R)$ gravity were studied. In this way, it becomes interesting to establish numerical relationships between the aforementioned shadow characteristics for Lovelock theories, which lack an analytical expression for their event horizon and are not asymptotically flat, but instead possess an AdS asymptote.
Regarding the aforementioned points, in this work, we will provide a methodology to study the analogy between gravitational tension and energy density in order to construct regular black hole solutions in LnFDGS gravity under the presence of a negative cosmological constant. It is worth mentioning that directly relating the gravitational tension to the Kretschmann scalar of the vacuum LnFDGS solution in the same way as for PL theory is complicated, because several additional terms would appear in the gravitational tension. Due to this, unlike Estrada and Aros (2025), we will define an alternative version of the Kretschmann scalar suitable for the LnFDGS AdS theory, and consequently, a redefinition of the gravitational tension. Additionally, we will speculatively discuss the connection between our model and the quantum Schwinger effect for this theory. We will test how the presence of the negative cosmological constant influences the obtained RBH solution, both in terms of its geometric and thermodynamic properties. Additionally, since it is not possible to find analytical relationships between the event horizon, the photon sphere radius, and the shadow size in LnFDGS, we will propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior.
II A brief revision of Lovelock gravity with $n$ fold degenerated ground state
Given that this family of theories is well known, it is worthwhile to recall only the relevant features for the discussion. First, the $d$-dimensional Lovelock (LL) Lagrangian is given by:
$L=\sum_{p=0}^{n=[d/2]}\frac{1}{2^{p}}\alpha_{p}\delta^{\nu_{1}\ldots\nu_{2}p}_{\mu_{1}\ldots\mu_{2}p}R^{\mu_{1}\mu_{2}}_{\hskip 6.02777pt\nu_{1}\nu_{2}}\ldots R^{\mu_{2p-1}\mu_{2p}}_{\hskip 6.02777pt\nu_{2p-1}\nu_{2p}}$ (2)
where $R^{\mu\alpha}_{\hskip 6.02777pt\nu\beta}$ is the Riemann tensor, $[X]$ stands for the integer part of $X$ and $\{\alpha_{p}\}$ is a set of dimensionful constants. The LL equations of motion can be written as
$\mathcal{G}^{\alpha}_{\hskip 3.01389pt\beta}=\sum_{p=0}^{n=[d/2]}\frac{d-2p}{2^{p}}\alpha_{p}\delta^{\alpha\nu_{1}\ldots\nu_{2}p}_{\beta\mu_{1}\ldots\mu_{2}p}R^{\mu_{1}\mu_{2}}_{\hskip 6.02777pt\nu_{1}\nu_{2}}\ldots R^{\mu_{2p-1}\mu_{2p}}_{\hskip 6.02777pt\nu_{2p-1}\nu_{2p}}.$ (3)
Here $\mathcal{G}^{(n)}_{\alpha\beta}$ denotes an $n$-th order generalization of the Einstein tensor, which is influenced by the order on the Riemann tensor of the Lagrangian $L$. For instance, $\mathcal{G}^{(1)}_{\alpha\beta}$ represents the Einstein tensor associated with the Ricci scalar (with the Einstein-Hilbert theory being a specific case of Lovelock theory), the $\mathcal{G}^{(2)}_{\alpha\beta}$ corresponds to the Lanczos tensor related to the Gauss-Bonnet Lagrangian and $\mathcal{G}^{(3)}_{\alpha\beta}$ corresponds to the cubic tensor.
Unfortunately, not even in (matter) vacuum these equations of motion can be solved straightforwardly for an arbitrary set of $\alpha$’s. Not even the ground-states, i.e. the locally constant curvature solutions, can be obtained in $d>8$ for a general set of $\alpha$’s. The problem of obtaining those constant curvatures solutions reduces to obtain the roots of a polynomial of order $[d/2]$, which, as it is well known, is not unattainable for $[d/2]>4$. Moreover, the roots can be complex numbers with nonvanishing imaginary parts. Therefore, to remove any problem, the asymptotic behavior is fixed to a single asymptotically locally flat or AdS behavior.
As mentioned in the introduction, Lovelock theory can be factorized in several effective cosmological constants Camanho and Edelstein (2013). In this connection, for $\alpha_{p}=0$ from $p>I$, the vacuum equations of motion can be written as Aros and Estrada (2019):
[…]
This shows, as expected, that the Lovelock gravity can be factorized in several ground states of constant curvature. In this work, we aim to complete the analysis considering asymptotically locally AdS spaces, i.e., spaces satisfying
$\left.R^{\mu\alpha}_{\hskip 6.02777pt\nu\beta}\right|_{x\rightarrow{\text{ALAdS region}}}\longrightarrow-\frac{1}{l^{2}}\delta^{\mu\alpha}_{\nu\beta}$ (5)
In the simplest cases where Eq. (5) is satisfied, the vacuum is $n$-fold degenerate, and the EOMs in Eqs. (3) and (4), after an appropriate choice of the coupling constants Crisostomo et al. (2000); Aros et al. (2001), take the simple form
[…]
with $0<n<[d/2]$.
The static vacuum solution of (6) was obtained in Crisostomo et al. (2000); Aros et al. (2001). In Schwarzschild coordinates, this is given by
$ds^{2}=-f(r)^{2}dt^{2}+\frac{1}{f(r)^{2}}dr^{2}+r^{2}d\Sigma_{\gamma}$ (7)
where $\Sigma_{\gamma}$ is a manifold of constant curvature $\gamma=\pm 1,0$, and where the vacuum solution is
$f(r)^{2}=\gamma+\frac{r^{2}}{l^{2}}-\left(\frac{2M}{r^{d-2n-1}}\right)^{\frac{1}{n}}.$ (8)
It must be stressed that $(X)^{1/n}$ stands for any of the $n$ different roots of (X). It is straightforward to check the presence of a singularity at $r=0$. Nonetheless, this is not enough, and in general, what characterizes the interesting physical solutions is that they satisfy $f(r)^{2}\in\mathbb{R}$, for $r>0$. If ∃$\exists\hskip 4.30554ptr^{*}$ such $f(r^{*})^{2}=0$ this defines a Killing horizon and thus the line element above Eqs.(7) and (8) describe a black hole geometry. One can notice how similar is this solution to the corresponding Schwarzschild solution  [^1] Nonetheless, at least there is significant difference. For $n$ even integer, since there are always two $\pm$ real roots in the expression, $M<0$ can be considered allowed..
It is worth stressing that, as mentioned in the introduction, the structure of the equations of motion leading to the last solution is quite different from that of PL, as PL considers only a single term in the Lovelock series, while LnFDGS includes all the terms of the mentioned series up to $n=[d/2]$ under the particular choice of coupling constants previously mentioned. This choice allows the equations of motion to be factorized into AdS vacua, as described above (which is not possible for PL).
III About the nature of the matter source proposed for Lovelock with $n$ fold degenerated ground state AdS.
Given the aforementioned differences between PL and LnFDGS theories, providing a form for the energy density analogous to the one described earlier is not straightforward. This is because several additional terms would appear when directly substituting the Kretschmann scalar of the vacuum LnFDGS solution. To address this, we will provide some ingredients. First, we will define gravitational tension through the gauge curvature tensor $F_{\mu\nu}$, which is the Poincaré (group) curvature. Second, we will define an appropriate tensor for LnFDGS, which we will call the Poincaré AdS-like curvature. From this, we will provide an alternative version of the Kretschmann scalar associated with gravitational tension. Finally, we will describe, from a physical perspective, the resulting structure of the energy density.
III.1 Defining gravitational tension through the gauge curvature and a glimpse of the analogy with the Schwinger effect.
Although the analogy between gravitational tension and the production ratio in the Schwinger effect requires further study, in this work we can speculate by providing some glimpses on the subject. As discussed in Ansoldi (2008); Dymnikova (1992), one could argue that the Schwinger effect should play a role in one of two ways. Firstly, it would give rise to corrections to the vacuum expectation value of the fields. These corrections can be expressed in terms of powers of the Riemann tensor Birrell and Davies (1982). Furthermore, gravitation itself should self-correct due to the presence of singularities. It is expected that these effects cannot be separated for large curvature. To address any of these approximations, after considering Eq. (1), one can recall that in asymptotically (Riemannian) flat spaces (i.e., with a vanishing cosmological constant), one analogy to the gauge curvature $F_{\mu\nu}$ is the Poincaré (group) curvature. In this way, we can also interpret the gravitational tension as $F=\sqrt{F_{\mu\nu}F^{\mu\nu}}$, where
$F^{\text{P}}_{\mu\nu}=e^{a}_{\hskip 3.01389pt\lambda}e^{b}_{\hskip 3.01389pt\rho}R^{\lambda\rho}_{\hskip 6.02777pt\mu\nu}J_{ab}+e^{a}_{\hskip 3.01389pt\rho}T^{\rho}_{\hskip 3.01389pt\mu\nu}P_{a}$ (9)
where $(J_{ab},P_{c})$ span the Poincare algebra. Here $R^{\lambda\rho}_{\hskip 6.02777pt\mu\nu}$ is the curvature tensor and $T^{\rho}_{\hskip 3.01389pt\mu\nu}$ the torsion tensor. However, since on a Riemannian manifold the torsion tensor vanishes and the curvature tensor is the Riemann tensor, the Schwinger effect could to be characterized by the Kretschmann scalar Ansoldi (2008); DYMNIKOVA (1996); Bena et al. (2021)
$F_{\mu\nu}F^{\mu\nu}\triangleq K=R^{\lambda\rho}_{\hskip 6.02777pt\mu\nu}R^{\mu\nu}_{\hskip 6.02777pt\lambda\rho}.$ (10)
Now, as mentioned earlier, in order to propose an energy density, one can observe that for the four-dimensional Schwarzschild solution, the Kretschmann scalar is proportional to $K\sim 1/r^{6}$, and thus, by substituting it into Eq. (1), the energy density proposed by Dymnikova is reproduced.
Before proceeding, it is worth mentioning that the prescription above has been further explored in many works, see Ansoldi (2008); DYMNIKOVA (1996); Estrada and Muniz (2023); Alencar et al. (2023). Essentially, the results are regular black hole and wormhole solutions. In Wondrak et al. (2023); Chernodub (2023), the mentioned relationship between the Kretschmann scalar and the Schwinger effect was explored in a different (gravitational) context.
III.2 An alternative definition of the Kretschmann scalar for our anti-de Sitter space.
To extend the framework to asymptotically (locally) Anti de Sitter spaces, it is enough to consider a $SO(d-1,2$) curvature Hassaine and Zanelli (2016) instead of the Poincare one. One option is Hassaine et al. (2005)
$F^{\text{AdS}}_{\mu\nu}=e^{a}_{\hskip 3.01389pt\lambda}e^{b}_{\hskip 3.01389pt\rho}\left(R^{\lambda\rho}_{\hskip 6.02777pt\mu\nu}+\frac{1}{l^{2}}\delta^{\lambda\rho}_{\mu\nu}\right)J_{ab}+e^{a}_{\hskip 3.01389pt\rho}T^{\rho}_{\hskip 3.01389pt\mu\nu}J_{a},$ (11)
In this work, for simplicity, we will refer to the tensor $F^{\text{AdS}}_{\mu\nu}$ as Poincaré AdS-like curvature, where $l$ is called the AdS radius and $J_{ab},J_{a}$ are a set of generators of $SO(d-1,2)$. Since the discussion is for Riemannian manifold, where $T^{\rho}_{\hskip 3.01389pt\mu\nu}=0$, in order to simplify the calculations, one can define an adequate Kretschmann scalar $K^{\prime}$
$(F_{\mu\nu})^{\text{AdS}}(F^{\mu\nu})^{\text{AdS}}\triangleq K^{\prime}=\left(R^{\lambda\rho}_{\hskip 6.02777pt\mu\nu}+\frac{1}{l^{2}}\delta^{\lambda\rho}_{\mu\nu}\right)\left(R_{\lambda\rho}^{\hskip 6.02777pt\mu\nu}+\frac{1}{l^{2}}\delta_{\lambda\rho}^{\mu\nu}\right)=K+\frac{8}{l^{2}}R+\frac{2}{l^{4}}(d-1).$ (12)
As done above, one can evaluate $K^{\prime}$ on the four-dimensional Schwarzschild-AdS solution, which yields
$K^{\prime}\sim\frac{1}{r^{6}}.$ (13)
This leads to the same energy density, see Eq.(1), of the $\Lambda=0$ case. This must be expected as for $r\approx 0$ the effects of the different asymptotical structure should be irrelevant.
III.3 The proposed model
As mentioned earlier, to define the energy density in (1), we will associate the gravitational tension with the value of our definition provided above for the Kretschmann scalar for the vacuum AdS LnFDGS solution. From equations (6) and (12)
$K^{\prime}=\left(\frac{d^{2}}{dr^{2}}g(r)\right)^{2}+\frac{2(d-2)}{r^{2}}\left(\frac{d}{dr}g(r)\right)^{2}+\frac{2(d-2)(d-3)}{r^{4}}\left(g(r)\right)^{2}$ (14)
where
$g(r)=\gamma+\frac{r^{2}}{l^{2}}-f(r)^{2}.$ (15)
Now, evaluating, using (8), yields
$K^{\prime}=\left(\frac{(d-n-1)^{2}}{n^{4}}+2(d-2)\left(\frac{1}{n^{2}}+(d-3)\right)\right)\frac{(2M)^{\frac{2}{n}}}{r^{\frac{2}{n}(d-1)}}$ (16)
Building on the idea previously described, where gravitational tension is proportional to the square root of the Kretschmann scalar of the vacuum solution
$F\sim\sqrt{K^{\prime}}\sim\frac{{M}^{\frac{1}{n}}}{r^{\frac{d-1}{n}}}$ (17)
In this way, the steps previously taken, namely: defining a tensor analogous to the gauge curvature, which we have called as PoincarĂŠ AdS-like curvature and providing an alternative definition for the Kretschmann scalar, now allow us to model an energy density. Thus, in analogy with (1), it can be defined the energy density
$\rho=A\exp\left(-\frac{r^{(d-1)/n}}{a^{(d-1)/n}}\right)$ (18)
where, for simplicity, the constant $A$ has been adjusted to:
$A=\frac{d-2}{n}\frac{M}{a^{d-1}/(d-1)}$ (19)
where $a$ is an arbitrary constant. It is worth mentioning that this density model satisfies what was described earlier. While the gravitational tension associated with the tidal forces of the vacuum LnFDGS AdS solution diverges near the origin, the energy density encodes the information of the aforementioned case in such a way that, at the origin, the density takes a finite value, which, as we will see below, is associated with the suppression of the usual singularity present at the radial origin.
IV A new static regular solution in Lovelock with $n$ fold degenerated ground state AdS
The energy-momentum tensor corresponds to a neutral perfect fluid:
$T^{\mu}_{\nu}=\mbox{diag}(-\rho,p_{r},p_{\theta},p_{\theta},...),$ (20)
On the one hand, it is well known that this form of the metric imposes the condition $\rho=-p_{r}$. To obtain the regular solution it is enough to solve (6)
$\mathcal{G}^{\alpha}_{\hskip 3.01389pt\beta}=T^{\alpha}_{\hskip 3.01389pt\beta}$ (21)
The components $(t,t)$ and $(r,r)$ of the equations of motion take the form:
$\frac{d}{dr}\left(r^{d-1}\left(\bar{f}_{\gamma}(r)+\frac{1}{l^{2}}\right)^{n}\right)=\frac{2}{d-2}r^{d-2}\rho,$ (22)
where $l$ corresponds to the AdS radius, which is related to the cosmological constant such that $\Lambda=-\frac{(d-1)(d-2)}{2l^{2}}$ and where
$\bar{f}_{\gamma}(r)=\frac{\gamma-f(r)^{2}}{r^{2}}$ (23)
On the other hand, due to the transverse symmetry, we have $p_{\theta}=p_{\phi}=p_{t}=\ldots$ for all the $(d-2)$ angular coordinates. Thus, using the aforementioned condition $\rho=-p_{r}$, the conservation law $T^{\alpha\beta}_{;\beta}=0$ gives:
$p_{t}=-\frac{r}{d-2}\rho^{\prime}-\rho$ (24)
Using the static ansatz (7) with
$f(r)^{2}=\gamma+\frac{r^{2}}{l^{2}}-\left(\frac{2m(r)}{r^{d-2n-1}}\right)^{\frac{1}{n}}$ (25)
the direct integration yields
[…]
By choosing $M\cdot(n-1)!$ as integration constant, the mass function is given by
$m(r)$ $=$ $M\left((n-1)!-\Gamma\left[n\,,\,\frac{r^{(d-1)/n}}{a^{(d-1)/n}}\right]\right)$ (27)
$=$ $\frac{M}{n}{}_{1}F_{1}\left(n,n+1,-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)\left(\frac{r}{a}\right)^{d-1}$
$=$ $M(n-1)!\left(1-\exp{\left(-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r}{a}\right)^{\frac{i}{n}(d-1)}\right)$
where ${}_{1}F_{1}$ is hyperbolic confluent function. See appendix A.
Finally, the solution can be written as
$f(r)^{2}$ $=$ $\gamma+\frac{r^{2}}{l^{2}}-\frac{r^{2}}{a^{2}}\left(\frac{2M}{n}{}_{1}F_{1}\left(n,n+1,-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)\right)^{\frac{1}{n}}$ (28)
$=$ $\gamma+\frac{r^{2}}{l^{2}}-\left(\frac{2M(n-1)!}{r^{d-2n-1}}\left(1-\exp{\left(-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r}{a}\right)^{\frac{i}{n}(d-1)}\right)\right)^{\frac{1}{n}}$
Before proceeding it is worthwhile noticing the coordinate $r$ defines the regions of interest. While $r\rightarrow\infty$ defines the asymptotic region, $r\rightarrow 0$ must define the center of the solution. Second, this solution has two parameters of distance competing, $l$ and $a$, and thus one can foresee that $a/l$ should have a relevant rôle in the analysis. Furthermore, given respective physical meaning of $a$ and $l$, in a physically sound scenario it must be satisfied $a/l\ll 1$. Because of that one can expect that if an event horizon exist, then the value of the radius of the external horizon, says $r_{+}$, should have only small deviation from one associated to the vacuum solution found in Aros et al. (2001).
One central feature of (28) is that any analysis of its behaviour can only be done numerically due to the hypergeometric function. This is quite forlon as in this case the numerical analysis bypasses the small corrections due to $a/l\ll 1$, the sound physical scenario. On the other hand, the silver lighting is that model simultaneously solves the singularities while proposing black hole geometries that are, in practice, indistinguishable from their (singular) counterparts.
IV.1 Behavior near $r=0$ region
Firstly one has to analyze the region near $r=0$, which basically separates the regular solution from its black hole counterpart. At $r=0$, one can show the lack of any singularity. This can be noticed by merely taking the limit
$\lim_{r\rightarrow 0}f(r)^{2}\rightarrow\gamma-\left(\left(\frac{2M}{n}\right)^{\frac{1}{n}}-\frac{a^{2}}{l^{2}}\right)\frac{r^{2}}{a^{2}}.$ (29)
In principle this defines either a (nearly) (Anti) de Sitter or a flat region nearby $r=0$. However, one can notice the presence of $a^{2}/l^{2}$ and thus, because in a realistic physical set-up $a/l\ll 1$, even for small masses $M$, recalled measured in term of units of $a^{d-3}$, it should be expected a de Sitter space region near $r=0$. This recovers the behavior without cosmological constant. Regarding the latter, in the following of this work, we will consider $\gamma=1$, i.e., the case where the transversal section is spherically symmetric.
From the last equation, it is straightforward to check that the Kretschmann invariant takes a finite value near the origin:
$\lim_{r\to 0}K=\frac{2(d^{2}-5d+8)}{a^{4}}\left(\frac{2M}{n}\right)^{2/n}$ (30)
IV.2 AdS asymptotic region
The firstly one needs to address how the regular solution in Eq.(28) evolves as the asymptotic region is approached, roughly as $r\rightarrow\infty$. As expected, the regular solution, as $r\rightarrow\infty$,
$\lim_{r\rightarrow\infty}f(r)^{2}\approx\gamma+\frac{r^{2}}{l^{2}}-\left(\frac{2M}{r^{d-2n-1}}\right)^{\frac{1}{n}}$ (31)
defining an asymptotically locally AdS space as $r$ grows but also a merging into the vacuum solution counterpart Crisostomo et al. (2000); Aros et al. (2001).
IV.3 Mass
Due to the AdS asymptotia of this family of solutions, the determination of the mass or energy is not straightforward, requiring a regularization process, see for instance Ashtekar and Das (2000); Aros et al. (2001); Skenderis (2002). Here, we use the methodology provided in reference Aros et al. (2001). In this case, given the simple static structure of the solution, any of those methods gives the same result
$E=M+E_{0}$ (32)
where $E_{0}$ vanishes in even dimensions and in odd dimensions stands for the vacuum energy of the AdS space Aros and Estrada (2019).
IV.4 Merging zone
One remarkable feature of the regular solution, see Eq.(28), is the existence, generically, of a value of the radial coordinate where the regular solution converges into the vacuum LnFDGS AdS solution found in references Crisostomo et al. (2000); Aros et al. (2001), which corresponds to Eq.(8).
Analytically, we note that the second factor inside the parentheses in equation (28) is such that:
$\lim_{r/a\rightarrow\infty}\left(1-\exp{\left(-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r}{a}\right)^{\frac{i}{n}(d-1)}\right)\approx 1,$ (33)
Which implies that for values such that $a\ll r\ll l$, our regular AdS solution and its vacuum AdS counterpart become almost indistinguishable.
However, analytically, it is difficult to estimate from which values of the radial coordinate our regular AdS solution and its vacuum AdS counterpart become indistinguishable. Numerically, we observe in Figure 1 that both cases can be indistinguishable for a finite value of $r_{-}<r=r_{*}<r_{+}$, greater than the inner horizon and less than the event horizon, or equivalently for values of $r$ where $f(r)^{2}<0$. Below, we will discuss some physical implications associated with this fact. It is worth emphasizing that $f(r)^{2}$ can have two, one, or zero roots depending on the values of $d,n,M,a$, and $l$. This is depicted generically in Fig.1.
Figure 1: This is how the merging occurs for different values of the mass parameter. Plot with $r$ on the horizontal axis and $f(r)^{2}$ on the vertical axis. The red curve corresponds to our regular solution, while the blue curve corresponds to the vacuum LnFDGS AdS solutions.
Even though this goes beyond the scope of this work, this rises up the obvious and general question of to what extend a regular solution, black hole or not, can be discerned from its vacuum black hole counterparts by an experiment carried out by an observer beyond that $r$ merging.
IV.5 Structure of horizons
To begin with the discussion it is worthwhile to stress that, and unlike vacuum LnFDGS AdS solutions in Crisostomo et al. (2000); Aros et al. (2001), due to the lack of singularities regular solutions does not need to be black holes, i.e. having a horizon, to be well defined. In other words, regular solutions are not affected by the cosmic censorship hypothesis.
Now rest to answer when a horizon is present for the static regular solution defined by (28). This occurs when $f(r)^{2}$ vanish, i.e., when the set of parameters $a,M$ and $l$ are such that (28) could vanish a given radius.
In figure 2, we can observe the generic behavior of the mass parameter versus radial values where the function $f^{2}(r)$ vanishes. The behavior in the vacuum LnFDGS AdS case Crisostomo et al. (2000); Aros et al. (2001) is illustrated in red, while the behavior of our RBH LUV AdS with our matter sources is shown in blue. Thus, the regions where $M$ is decreasing correspond to the values of the inner horizon, and the regions where $M$ is increasing correspond to the values of the black hole horizon in our solution. Therefore, for $M>M_{min}$, there exists an inner horizon $r_{-}$ and a black hole horizon $r_{+}$ for the same value of the parameter $M$. The value $M_{min}$ corresponds to where the inner and black hole horizons coincide, i.e., an extremal black hole. A notable feature is that for a small deviation to the right in the value of the extremal radius $r_{ext}=r_{-}=r_{+}$ , i.e $r_{*}>r_{ext}$ (which coincides with a small upper deviation of the value of $M$ with respect to $M_{min}$), our regular AdS solution and its vacuum AdS counterpart become indistinguishable, as the red and blue curves are also indistinguishable.
[…]
As a consequence of the above, we also note that for values considerably larger than the Planck scales, such as the observed values of mass and radius of rotating black holes observed in LIGO GW150914 Abbott et al. (2017), both the vacuum and regular solutions are also indistinguishable. Thus, it becomes intriguing to observe the properties of black holes at scales much smaller than those that have been observed to date.
Figure 2: For single value of $M$ there are an interior and exterior horizons. Furthermore, one notice that large exterior $r_{+}$ the vacuum solution is approached
V Shadows
We study the line element in (7) for the spherically symmetric case with $\gamma=1$. Thus, the transversal section can be written as:
$d\Omega_{D-2}=d\theta^{2}_{1}+\sum_{j=2}^{D-2}d\theta^{2}_{j}\left(\prod_{k=1}^{j-1}\sin^{2}\theta_{k}\right)$ (34)
We can notice that the metric tensor is independent of the angular coordinate $\theta_{D-2}=\phi$ and the temporal coordinate. In this way, we can identify the following Killing vectors:
$K_{\mu}=\left(-f(r),0,0,0,...\right)$ (35)
$K_{\mu}=\left(0,0,0,...,r^{2}\left(\prod_{k=1}^{D-3}\sin^{2}\theta_{k}\right)\right)$ (36)
These Killing vectors lead to conserved quantities such that:
$K_{\mu}\dot{x}^{\mu}=\mbox{constant}$ (37)
where the dot indicates the derivative respect to the affine parameter. We consider a movement of a photon in an equatorial plane described by $\theta_{1}=\theta_{2}=\ldots=\theta_{D-3}=\pi/2$ and $\theta_{D-3}$ constant. Thus
$f(r)\dot{t}=E$ (38)
$r^{2}\dot{\phi}=L$ (39)
where, in our case, $E$ and $L$ will represent the energy and angular momentum of the photon as measured by an observer at infinity.
It is straightforward to check that the norm of the tangent vector to the geodesic is also conserved. In this line, we have:
$\epsilon=-g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}$ (40)
Use $\epsilon=0$ for null geodesics. For time-like geodesics (massive particles), $\epsilon=1$, and for space-like geodesics, $\epsilon=-1$. As indicated in Paithankar and Kolekar (2023), due to the spherical symmetry, it is sufficient to consider the motion of null geodesics in an equatorial plane. From the previous equations, we obtain:
$\frac{\dot{r}^{2}}{L^{2}}+V_{eff}(r)=\frac{1}{b^{2}}$ (41)
where the effective potential is:
$V_{eff}(r)=\frac{f(r)}{r^{2}}$ (42)
and where the impact parameter $b$ is the perpendicular distance between the position of a photon moving towards the black hole and its center.
We say that the radius of the circular orbits is the distance between the center of the black hole and a point in spacetime where photons follow a circular motion. The geodesic for these distances is a closed circle. This forms the so-called photon sphere, whose radius we will denote as $r_{sp}$.
Minima in the effective potential function correspond to stable circular orbits. Maxima, on the other hand, correspond to unstable circular orbits. It is important to mention that, in several solutions such as the Schwarzschild solution or the vacuum black hole solution of Pure Lovelock, the photon sphere is unstable because it corresponds to a maximum of the effective potential Paithankar and Kolekar (2023). However, it plays an important physical role. The photon sphere is the union of all closed null geodesics. Thus, there are no other types of closed null geodesics in the spacetime. Photons tangential to the photon sphere will remain on the photon sphere. However, since the photon sphere is unstable, small perturbations will cause the photon to either fall into the black hole or escape from it. In this way, the radius of the photon sphere is determined by the following equation:
$00=\frac{dV_{eff}}{dr}\bigg{|}_{r=r_{sp}}=r_{sp}^{-3}\left(r_{sp}f^{\prime}(r)|_{r=r_{sp}}-2f(r_{sp})\right)\Rightarrow r_{sp}f^{\prime}(r)|_{r=r_{sp}}-2f(r_{sp})=$ 0 (43)
wherethe prime indicates differentiation with respect to the radial coordinate. Following reference Paithankar and Kolekar (2023), for simplicity, we assume that all light sources are placed uniformly at infinity. The critical impact parameter of the system $b_{cr}$ is the impact parameter for which the photon will precisely fall into the circular orbit around the black hole. For an impact parameter smaller than the critical impact parameter, the photon will always end up inside the black hole. For an impact parameter larger than the critical impact parameter, the photon will end up at an infinite distance from the black hole. Following Paithankar and Kolekar (2023), an observer located at infinity leads to the critical impact parameter being equal to the shadow radius, $r_{sh}=b_{cr}$. See the scheme in Figure 3. This can be computed by setting $\dot{r}=0,r=r_{sp}$ in equation (43):
$r_{sh}=b_{cr}=\frac{1}{\sqrt{V_{eff}}}=\frac{2\sqrt{f(r_{sp})}}{f^{\prime}(r)|_{r=r_{sp}}}$ (44)
Figure 3: Scheme
Given the lack of an analytical expression for the event horizon, the photon sphere radius, and the shadow size in our case study, and since our case is not asymptotically flat but instead possesses an AdS asymptote, it becomes interesting to establish numerical relationships between the aforementioned shadow characteristics for our case study of Lovelock theories. In the subsection below, we propose a method to achieve this. We will use the case $n=2,d=7$ as an example, but it is straightforward to check that our proposal is applicable to other values of $n$ and $d$.
V.1 A recipe for determining the size of the photon sphere and the shadow.
  1. 1.
    First, since it is not possible to obtain an analytical value for the photon sphere radius from equation (43), we obtain the mass parameter from this equation for our RBH solution, equations (25) and (27).
    $M(r_{sp})=$
    $\frac{2^{1-\frac{1}{n}}e^{\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}}n^{2}\left((-1+n)!-\Gamma\left[n,\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}\right]\right)r_{sp}^{(1-d)}\left((-1+n)!-\Gamma\left[n,\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}\right]\right)^{-\frac{1}{n}}}{(-1+d)r_{sp}^{2}\left(-\left(\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}\right)^{n}+e^{\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}}n(-1+n)!-e^{\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}}n\Gamma\left[n,\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}\right]\right)}$ (45)
    As an example, we will study the case $n=2,d=7$ in figure 4. The values of the ascending curve (In the example, as we can see in the orange dashed vertical line, values greater than ∼$\sim 6.5$ on the horizontal axis) represent the candidate values of the photon sphere radius. We use the word ”candidate” because later we need to ensure that the condition $r_{sp}>r_{+}$ is satisfied. The vertical axis shows the values of the mass parameter. Thus, each ordered pair $\left(r_{sp},M(r_{sp})\right)$ represents values of the candidate photon sphere and the mass parameter in the parameter space, such that equation (43) is satisfied, i.e., where the potential reaches an unstable maximum. On the other hand, the orange dashed horizontal line in figure 4 represents the maximum value of the mass parameter for the parameters used. Below, we will detail what the black horizontal line in the figure represents.
    Figure 4: In the horizontal axis the candidate values of the photon sphere radius $r_{sp}$. In the vertical axis the mass parameter $M(r_{sp})$. Using $n=2,d=7,a=4$
  2. 2.
    From equation (44) we can obtain the following relation between the photon sphere radius and the shadow’s radius:
    $r_{sh}=\frac{r_{sp}}{\sqrt{1+\frac{r_{sp}^{2}}{l^{2}}+\frac{2e^{\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}}n^{2}\left((-1+n)!-\Gamma\left[n,\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}\right]\right)}{(-1+d)\left(\left(\frac{r_{sp}}{a}\right)^{(d-1)}-e^{\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}}n(-1+n)!+e^{\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}}n\Gamma\left[n,\left(\frac{r_{sp}}{a}\right)^{\frac{d-1}{n}}\right]\right)}}}$ (46)
    Since equation (1) is independent of the parameter $l$, we can numerically test the values of this parameter where the condition that the shadow radius is greater than the photon sphere radius is satisfied. To do this, we construct the following function in the parameter space:
    $g(r_{sp},l)=\begin{cases}1&\text{if }r_{sh}>r_{sp},\\ 0&\text{if }r_{sh}<r_{sp}.\end{cases}$ (47)
    In this way, we observe in figure 5 that, for the parameters used, the condition $r_{sch}>r_{sp}$ is satisfied for approximately $l>41$ for all value of $r_{sp}$. Therefore, in the next step, we must use a value of $l$ that satisfies the aforementioned condition.
    Figure 5: $g(r_{ph},l)$ using $n=2,d=7,a=4$
  3. 3.
    For simplicity, we will write the mass parameter $M(r_{+})$, which depends on the event horizon $r_{+}$, in the parameter space as:
    $M(r_{+})=\frac{r_{+}^{-1+d-2n}\left(1+\frac{r_{+}^{2}}{l^{2}}\right)^{n}}{2\left((-1+n)!-\Gamma\left[n,\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right]\right)}$ (48)
    In figure 6, we display the behavior of the mass parameter. We must be consistent with the previous point and choose a value of $l$ that satisfies what was described there. In this way, on the ascending curve, we have an ordered pair of the form $\left(r_{+},M(r_{+})\right)$. The black dashed horizontal line represents the extremal minimum value of the mass for the chosen parameters, emphasizing that $l$ is not arbitrary.
    Figure 6: The values on the horizontal axis of the ascending curve represent the event horizon radius $r_{+}$. On the vertical axis, we have the mass parameter $M(r_{+})$. Using $n=2,d=7,a=4,l=100$.
  4. 4.
    In Figure 4, we have ordered pairs $\left(r_{sp},M(r_{sp})\right)$, while in Figure 6, we have the ordered pairs $\left(r_{+},M(r_{+})\right)$. In this regard, we note the following:
      The mass parameter must be the same in both cases when the remaining parameters are fixed. Thus, to numerically test the relationship between $r_{+}$ and $r_{sp}$, we must consider that
      $M(r_{sp})=M(r_{+})$ (49)
      Following this condition, the minimum value of the mass parameter is given by the extremal value, represented by the black dashed horizontal line in Figure 6, $M(r_{+})_{min}=M(r_{ext})$ (in the example, ∼$\sim 20$). Therefore, the lower bound for the photon sphere radius is given by $r_{sp}$ such that $M(r_{sp})=M(r_{+})_{min}=M(r_{ext})$ (vertical line in Figure 4, where $r_{sp}\sim 9.8$ and $M(r_{sp}\sim 9.8)\sim 20$). On the other hand, in the orange dashed line of Figure 4, it can be seen that the parameter $M(r_{sp})$ has a maximum value. This last value sets the upper bound for the photon sphere and event horizon, such that condition (49) is satisfied ($r_{sp}\sim 36$ and $r_{+}\sim 23.5$ in the examples of Figures 4 and 6, respectively).
      Once the appropriate ranges for $r_{+}$ and $r_{sp}$ have been determined in the previous steps, we plot the relationship between $r_{+}$ and $r_{sp}$ in Figure 7. This has been done numerically, using condition (49) on the ordered pairs $\left(r_{sp},M(r_{sp})\right)$ and $\left(r_{+},M(r_{+})\right)$ from Figures 4 and 6, respectively. As expected, we observe that the value of the photon sphere is greater than that of the event horizon. That is, photons follow a circular motion due to the geometric distortion of spacetime caused by the presence of the black hole.
      Figure 7: The values on the horizontal axis represent the event horizon radius $r_{+}$. The values on the vertical axis represent the sphere photon radius $r_{sp}$. . Using $n=2,d=7,a=4,l=100$.
  5. 5.
    Once the appropriate ranges for $r_{sp}$ and $l$ have been determined in the previous steps, using equation (46), we plot the relationship between the photon sphere radius and the shadow radius in Figure 8. As expected, we observe that the value of the photon sphere radius is smaller than that of the shadow radius. That is, photons that are directed with an impact parameter equal to $r_{sh}$ towards the black hole curve their trajectory towards the unstable photon sphere.
    Figure 8: The values on the horizontal axis represent the event horizon radius $r_{+}$. The values on the vertical axis represent the sphere photon radius $r_{sp}$. . Using $n=2,d=7,a=4,l=100$.
Following the methodology described earlier, in Figure 9 we have displayed the behavior of the event horizon versus the photon sphere radius for $n=2,d=6,7,8$ (left panel) and $n=3,d=8,9,10$ (right panel). It is important to note that, in agreement with the analysis carried out above, for each case, depending on the values of $n,d,l,a$ in the parameter space, there are different ranges for the values of the event horizon and the photon sphere radius. We can note that, if we use a common event horizon value as a reference for the same value of $n$, the photon sphere radius decreases as the number of dimensions increases.
Figure 9: The values on the horizontal axis represent the event horizon radius $r_{+}$. The values on the vertical axis represent the sphere photon radius $r_{sp}$. In left panel we have used $n=2,a=4,l=100$, $d=6$(red), $d=7$(blue), $d=8$(black). In right panel we have used $n=3,a=4,l=100$, $d=8$(red), $d=9$(blue), $d=10$(black)
In Figure 10, we have displayed the behavior of the photon sphere radius versus the black hole shadow radius for $n=2,d=6,7,8$ (left panel) and $n=3,d=8,9,10$ (right panel). In the same way as in the previous figure, and in agreement with the analysis carried out above, for each case, depending on the values of $n,d,l,a$ in the parameter space, there are different ranges for the values of the photon sphere radius and the shadow radius. We can note that, if we use a common photon sphere radius value as a reference for the same value of $n$, the size of the black hole shadow radius tends to decrease as the number of dimensions increases.
It is also important to mention that in all the cases studied numerically, the condition $r_{+}<r_{sp}<r_{sh}$ is satisfied.
These results could eventually be used for a deeper study in a future paper, estimating, for example, the deflection angle or the shape of a linearly uniformly accelerated trajectory, which can be interpreted as the curved spacetime generalization of the hyperbolic Rindler trajectory in flat spacetime. Both topics are beyond the scope of this work.
Figure 10: The values on the horizontal axis represent the sphere photon radius $r_{sp}$. The values on the vertical axis represent the shadow radius $r_{sh}$. In left panel we have used $n=2,a=4,l=100$, $d=6$(red), $d=7$(blue), $d=8$(black). In right panel we have used $n=3,a=4,l=100$, $d=8$(red), $d=9$(blue), $d=10$(black)
VI Thermodynamics
From now on, only the black hole solutions, the case where $M>M_{min}$, will be studied. To study the thermodynamics is helpful to express the mass, $M$, in terms of $r_{+}$ as follows
$M(r_{+})=n\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)^{n}r_{+}^{d-2n-1}\left(1-\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)^{-1}$ (50)
Analytically, we note that for $r_{+}/a\rightarrow\infty$, this mass function converges to the vacuum solution. However, as mentioned above, from Figure 2 we can observe numerically that for a value of the event horizon smaller than the one mentioned, which is slightly greater than the extremal radius, $r_{*}>r_{ext}$, the mass parameter of both cases becomes indistinguishable.
VI.1 Temperature
The temperature is defined by
$4\pi T$ $=$ $\frac{d}{dr}f(r)^{2}=2\frac{r_{+}}{l^{2}}+\frac{(d-2n-1)}{nr_{+}}\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)$ (51)
-- $\frac{1}{n!}\frac{\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}}{\left(1-e^{-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)}\left(\frac{r_{+}}{a}\right)^{\frac{n-1}{n}(d-2n-1)}$
To analyze the expression above it is worth to recall the presence of the two length scales $a$ and $l$. One can notice that the two first terms of Eq.(51) not only are independent of $a$, but coincide with the expression of the temperature found in Aros et al. (2001); Crisostomo et al. (2000) for the vacuum LnFDGS AdS solution. Therefore, to address the difference, one must play attention to the last term. In this connection, one can notice that this term dependents only on $r_{+}/a$. The structure of the third term for $r_{+}\gg a$ has the form
$\lim_{r_{+}/a\rightarrow\infty}\frac{\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}}{\left(1-e^{-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)}\left(\frac{r_{+}}{a}\right)^{\frac{n-1}{n}(d-2n-1)}\approx 0,$ (52)
Thus, we analytically observe that in a scenario where $a\ll r_{+}$, the temperature of the vacuum LnFDGS AdS solution becomes indistinguishable from our RBH LnFDGS AdS solution.
Numerically, we can note that for the value of $M=M_{min}$ where $r_{-}=r_{+}=r_{ext}$, the temperature vanishes. In this sense, it is well known that the zero temperature point is associated with a black remnant, known as what remains once evaporation stops. In this context, in Figure 11 below, we observe that a remnant indeed forms in this extremal case.
[…]
In other words, the behavior of temperature suggests that our forms of matter in the energy-momentum tensor play a role in eliminating divergences in LnFDGS AdS theories. This is analogous to the role of non-commutative matter sources in General Relativity (GR), which in turn play the same role in quantum field theory and string theory Nicolini et al. (2006); Smailagic and Spallucci (2013) . Therefore, it would be of interest to study the role of our LnFDGS AdS energy sources in quantum field theory in a future work. As shown, the role of our matter sources is to cool the black hole in the final stage. Consequently, analogous to the non-commutative $4D$ case studied in GR Nicolini et al. (2006), this could be interpreted as a suppression of the quantum back-reaction of Hawking radiation once the temperature has already reached its maximum, as the black hole emits progressively less energy. As mentioned, these effects could be tested at Planck scales.
From Figure 2, we can check that the sign of the derivative $\frac{dM}{dr_{+}}$ is always positive, i.e., $\frac{dM}{dr_{+}}\geq 0$. Thus, the sign of the derivative $\frac{dT}{dM}=\left(\frac{dT}{dr_{+}}\right)\left(\frac{dM}{dr_{+}}\right)^{-1}$ depends only on the sign of the derivative $\frac{dT}{dr_{+}}$. In summary, for our solution, the temperature exhibits three regions from left to right: i) $\frac{dT}{dM}>0$ due to the action of matter sources at small scales, ii) and iii) which coincide with the vacuum LnFDGS AdS case, where $\frac{dT}{dM}<0$ and $\frac{dT}{dM}>0$, respectively.
Figure 11: The temperature for $d=6$, $n=2$, $a=1$ and $l=10$
VI.2 Heat Capacity
Given that there is just a single parameter in this case, says $r_{+}$, is the heat capacity can be written as
$C=\frac{dE}{dT}=\left(\frac{dT}{dM}\right)^{-1}=\frac{dM}{dr_{+}}\left(\frac{dT}{dr_{+}}\right)^{-1}$ (53)
As mentioned before, from Figure 2, the derivative $\frac{dM}{dr_{+}}$ is always positive, i.e., $\frac{dM}{dr_{+}}\geq 0$. Since the sign of the derivative $\frac{dT}{dM}=\left(\frac{dT}{dr_{+}}\right)\left(\frac{dM}{dr_{+}}\right)^{-1}$ depends only on the sign of the derivative $\frac{dT}{dr_{+}}$, consequently, the sign of the heat capacity also depends solely on this last derivative.
Before to compute the heat capacity, it is worth to recall the heat capacity of the vacuum LnFDGS AdS case Crisostomo et al. (2000); Aros et al. (2001). This is given by
$C_{\text{vac}}=\frac{\left(\frac{\gamma(d-2n-1)}{2r_{+}}+(d-1)\frac{r_{+}}{2l^{2}}\right)\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)^{n-1}r_{+}^{d-2n-1}}{\left(\frac{2}{l^{2}}-\frac{(d-2n-1)}{nr_{+}^{2}}\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)+\frac{2(d-2n-1)}{nl^{2}}\right)}$ (54)
One can notice $C_{\text{vac}}$ has a divergency. This defines the phase transition noticed in Crisostomo et al. (2000).
To compute the heat capacity in this case one only need to consider the mass function Eq.(50) and the temperature (51) above.
$C$ $=$ $\frac{\left(\frac{\gamma(d-2n-1)}{2r_{+}}+(d-1)\frac{r_{+}}{2l^{2}}\right)\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)^{n-1}r_{+}^{d-2n-1}}{\left(\frac{2}{l^{2}}-\frac{(d-2n-1)}{nr_{+}^{2}}\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)+\frac{2(d-2n-1)}{nl^{2}}\right)}\left(1-\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)^{-1}$ (55)
-- $\frac{n\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)^{n}r_{+}^{d-2n-1}}{\left(\frac{2}{l^{2}}-\frac{(d-2n-1)}{nr^{2}}\left(\gamma+\frac{r^{2}}{l^{2}}\right)+\frac{2(d-2n-1)}{nl^{2}}\right)}\times n\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}\left(\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)^{n-1}$
$\left(1-\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)^{-2}\left(\frac{d-1}{n}\right)\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}-1}$
This, in turn, allows one to analytically conclude that
$\lim_{r_{+}/a\rightarrow\infty}C=C_{\text{vac}}$ (56)
This is quite interesting in two aspects. Firstly, due to $r_{+}/a\gg 1$, as mentioned earlier, it is a natural consideration for solar mass black holes, and therefore the standard thermodynamics must be recovered in that regime. Secondly, it must be recalled that, unlike for non-regular black holes, the radius of the outer horizon is bounded by $r_{ext}<r_{+}$ and thus cannot be arbitrarily small. In this way, the limit of interest is $r_{ext}\rightarrow r_{+}$, which represents the limit as $T\rightarrow 0$.
VI.2.1 On the radial evolution
A usual interpretation of the sign of the heat capacity is as follows Estrada and Aros (2023, 2025): a positive heat capacity indicates that when the temperature decreases (increases), the black hole emits (absorbs) thermal energy, thus $dM<0$ ($dM>0$) in the black hole, in order to reach thermodynamic equilibrium with the external environment, i.e., the black hole is stable. Conversely, a negative heat capacity implies that, if the temperature increases (decreases), the black hole also emits (absorbs) thermal energy toward the external environment, i.e., the black hole is unstable. As we will see below, under this interpretation, we can assume that at Planck scales, our matter sources induce a phase transition such that the black hole becomes stable at these scales, leading to a black hole remnant rather than a divergent temperature.
On the other hand, related to the analysis of phase transitions, this can be easily unfolded by using the fact that $C=\left(\frac{dT}{dM}\right)^{-1}\sim\left(\frac{dT}{dr_{+}}\right)^{-1}$ in graph 11. Moving from right to left, we can notice that once the vacuum LnFDGS AdS solution and our RBH LnFDGS AdS solution are no longer indistinguishable, the heat capacities of both solutions also differ. While the heat capacity of the vacuum AdS solution remains negative, meaning that its complete evaporation, as the parameter $M$ decreases to $M\to 0$, would imply the temperature having to rise to infinity, in our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that, as the mass decreases to reach the extremal value $M=M_{\text{min}}$, the temperature decreases until the black hole remnant is reached at $T=0$, where the evaporation process halts. As mentioned earlier, this could occur at scales close to the Planck scale. That is, the nature of our matter sources, along with avoiding the formation of a singularity, would lead to the formation of a remnant. Thus, in our solution, there would be two phase transitions, from right to left: i) $C>0$ where $dT/dM>0$, ii) $C<0$ where $dT/dM<0$ (these two coincide with the vacuum LnFDGS AdS case, where both solutions are indistinguishable), iii) $C>0$ where $dT/dM>0$.
VII DISCUSSION AND SUMMARIZE
Motivated by the following facts:
    The recent higher-dimensional model to suppress physical singularities from Estrada and Aros (2025) for Pure Lovelock gravity, where the energy density encodes the gravitational information of the vacuum solution through the Kretschmann scalar.
    The theoretical importance of incorporating a negative cosmological constant in various physical problems such as the AdS/CFT correspondence, the thermodynamic interpretation of the negative cosmological constant as a positive pressure in the universe, etc.
    The invalidity of including a negative cosmological constant in PL theories in the spherically symmetric case, since its inclusion leads to the appearance of a curvature singularity. This singularity is such that, although the metric tensor is regular for a radius greater than the event horizon, the Ricci and Kretschmann invariants diverge at a point within this location.
    It is worth mentioning that directly relating the gravitational tension to the Kretschmann scalar of the vacuum LnFDGS solution in the same way as for PL theory is complicated, because several additional terms would appear in the gravitational tension.
    Related to the study of shadows, recently in reference Paithankar and Kolekar (2023), relationships were found between the event horizon, the photon sphere radius, and the shadow size for the vacuum solution without a cosmological constant in Pure Lovelock. However, the inclusion of matter and the cosmological constant makes it impossible to find analytical relationships for our case study of LnFDGS AdS. Therefore, it is useful to propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior.
To address the points described above, instead of considering the PL action, which takes into account only one term in the Lovelock series, we use the definition of the coupling constants from references Crisostomo et al. (2000); Aros et al. (2001), such that, by considering all the terms in the Lovelock series up to order $n=[d/2]$, the equations of motion possess an $n$-fold degenerate ground state AdS. In this way, the inclusion of the negative cosmological constant does not lead to the appearance of a curvature singularity.
Given the structure of the equations of motion in LnFDGS theories, providing a form for the energy density analogous to the one described earlier is not straightforward. This is because several additional terms would appear when directly substituting the Kretschmann scalar of the vacuum LnFDGS solution. To address this, we have defined an appropriate gauge curvature tensor for LnFDGS, $F^{\text{AdS}}_{\mu\nu}$, which we have called the Poincaré AdS-like curvature. From this, we will provide an alternative version of the Kretschmann scalar associated with gravitational tension. In this way, using the structure provided for gravitational tension, we have defined the energy density, which maintains the physical arguments provided in reference Estrada and Muniz (2023) for PL.
The obtained solution is such that there exists a value of the radial coordinate $r_{-}<r=r_{*}<r_{+}$, greater than the inner horizon and less than the event horizon, such that for radial values slightly greater than $r_{*}$, the vacuum and singular LnFDGS solution from references Crisostomo et al. (2000); Aros et al. (2001) becomes indistinguishable from our regular LnFDGS solution. In the case that $r_{*}$ is of the order of the Planck scales, the geometric differences between both solutions, and specifically, the suppression of the singularity, would occur at quantum scales.
Related to the study of shadows, a method has been proposed to obtain numerical graphical relationships between the event horizon, the photon sphere radius, and the shadow size in our case study. The case $n=2,d=7$ has been used as an example to explain our procedure. However, it is straightforward to check that our proposal is applicable to other values of $n$ and $d$. We have also displayed the graphical behavior for the cases $n=2,d=6,7,8$ and $n=3,d=8,9,10$. We can note that, if we use a common event horizon value as a reference for the same value of $n$, the photon sphere radius decreases as the number of dimensions increases. Additionally, if we use a common photon sphere radius value as a reference for the same value of $n$, the size of the black hole shadow radius tends to decrease as the number of dimensions increases. It is also important to mention that in all the cases studied numerically, the condition $r_{+}<r_{sp}<r_{sh}$ is satisfied.
From the analysis of the temperature, it can be observed that there exists a finite value slightly greater than the extremal radius, $r_{*}>r_{ext}$, where the temperatures of both cases also become indistinguishable. Thus, at the Planck scales, quantum effects would arise, meaning that instead of the temperature evolving to infinity as in the vacuum LnFDGS case, the matter sources proposed in this work cause the temperature to decrease until a black hole remnant is reached at $T=0$ and $r_{+}=r_{ext}$.
Regarding the analysis of heat capacity and radial evolution, moving from right to left in graph 11, we can notice that once the vacuum LnFDGS AdS solution and our RBH LnFDGS AdS solution are no longer indistinguishable, the heat capacities of both solutions also differ. While the heat capacity of the vacuum AdS solution remains negative, meaning that its complete evaporation, as the parameter $M$ decreases to $M\to 0$, would imply the temperature rising to infinity, in our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that, as the mass decreases to reach the extremal value $M=M_{\text{min}}$, the temperature decreases until the black hole remnant is reached at $T=0$, where the evaporation process halts. We could assume that at Planck scales, our matter sources induce a phase transition such that the black hole becomes stable at these scales, leading to a black hole remnant rather than a divergent temperature.
appendices
Appendix A Hypergeometric Confluent Function Expression
The confluent hypergeometric function above can be written as
${}_{1}F_{1}\left(n,n+1,-y\right)=\frac{n!}{y^{n}}\left(1-e^{-y}\sum_{i=0}^{n-1}\frac{y^{i}}{i!}\right)$ (57)
and thus in the case at hand
${}_{1}F_{1}\left(n,n+1,-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)=n!\left(\frac{a}{r}\right)^{d-1}\left(1-\exp{\left(-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r}{a}\right)^{\frac{i}{n}(d-1)}\right)$ (58)
The derivative of this function is given by
$\frac{d}{dy}\left({}_{1}F_{1}\left(n,n+1,-y\right)\right)=\frac{n}{y}\left(e^{-y}-{}_{1}F_{1}\left(n,n+1,-y\right)\right)$ (59)
Acknowledgements
This work of RA was partially funded through FONDECYT-Chile 1220335. Milko Estrada is funded by the FONDECYT IniciaciĂłn Grant 11230247.